Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
More Details for 2007-11-22
ISS On-Orbit Status 11/22/07

Happy Thanksgiving on Earth and in Heaven (and in between)!

CDR Whitson & FE-1 Malenchenko started off on today's light-duty schedule with another standard 30-min Shuttle RPM (R-bar Pitch Maneuver) skill training, Peggy's third, Yuri's fourth, using DCS-760 digital still cameras with 400 & 800mm lenses at Service Module (SM) windows 6 & 8 to take imagery of documented EO (Earth Observation) targets facing the velocity vector (in flight direction). Afterwards, Peggy downlinked the obtained images to the ground for analysis, to be discussed at a subsequent tagup. (The skill training prepares crewmembers for the bottomside mapping of the Orbiter at the arrival of STS-122/1E in December. During the RPM at ~600 ft from the station, the ISS crew will have only ~90 seconds for taking high-resolution digital photographs of all tile areas and door seals on the Atlantis from SM windows 6 & 8, to be downlinked for launch debris assessment. Thus, time available for the shooting will be very limited, requiring great coordination between the two headset-equipped photographers and the Shuttle.)

Afterwards, the FE-1 continued the current round of periodic preventive maintenance of cabin ventilation systems in the RS (Russian Segment), today cleaning 'Group A' fan grilles in the SM.

The CDR conducted the weekly 10-min. CWC (Contingency Water Container) audit as part of the continuing WDS (Water Delivery System) assessment of onboard water supplies. (Updated 'cue cards' based on the crew's water calldowns are sent up every other week. Analysis of samples of water transferred in CWCs from Discovery to ISS during 10A confirmed bacterial contamination in all containers transferred. Currently off-limit for crew use until further word are two CWCs with potable water and nine CWCs with technical water. Background: The identified contaminant, a common soil bacterium (unicellular organism) called Wautersia after Belgian microbiologist Georges Wauters, is no more critical than what is found often in faucet water on the ground or in farm soil. Wautersia lives off hydrogen & carbon dioxide, oxidizing H2 and producing oxyhydrogen as energy for itself. Since it can turn sugar into a synthetic biodegradable fuel, it was seen for a short while as a promising long-term solution to the petroleum dependency, until it became clear that it would require gigantic amounts of expensive sugar.)

The FE-2 performed his fourth ICEPAC insertion in the MELFI (Minus Eighty-Degree Laboratory Freezer for ISS), retrieving two more -32 degC ICEPAC belts from stowage and placing them in Dewar 2/Tray B/Section 1 & Tray A/Section 4. (The reason the crew is performing several ICEPAC insertions this week is because the amount of warm mass that can be placed in a dewar at one time is limited. These activities are in preparation for the next Cold Bag packing, planned for STS-122/1E.)

In the SM, Malenchenko conducted his first periodic repositioning of the ESA/RSC-Energia experiment ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS) by changing the position of its AST spectrometer on panel 437 (90 deg rotation in its place to face in the Z- or upward direction) and swapping the AST's PCMCIA (Portable Computer Memory Card International Adapter) storage card #936 with #937. The activity was photo-recorded, with imagery downlinked afterwards. (ALTCRISS uses the AST spectrometer to monitor space radiation in the Russian segment (RS).)

Continuing their preparations for the next spacewalk, EVA-12 'Charlie' on 11/24 (Saturday), EV1 Peggy Whitson & EV2 Dan Tani worked in the U.S. Airlock (A/L) where they -

Readied EVA tools (including checking a tether for EV2),
Recharged the EMU backpacks with water (EMU #3006 got a full water tank dump & recharge; #3018 had only its water supply recharged),
Reviewed uplinked updated procedural material for the spacewalk (some tasks, including get-aheads, were added since yesterday, see below),
Terminated regeneration of METOX (Metal Oxide) canister #0015 & #0019 in the A/L bakeout oven and started the process on canisters #0020 & #0021 (the A/L, currently Dan Tani's sleeping quarters, is staying within allowable ppCO2 (Carbon Dioxide Partial Pressure) limits for Dan's sleeptime), and
Terminated BSA (Battery Stowage Assembly) recharging of the EMU batteries used in EVA 'Bravo'.
(Background: The regenerable METOX absorbers of CO2 are currently not used in the EMUs but in the A/L during tomorrow night's 'Campout' lockout of Peggy & Dan. During their spacewalk, CO2 scrubbing in the EMUs will be handled by LiOH (Lithium Hydroxide) canisters, with more 'lifetime' than METOX cans, to protect against possible timeline extension of EVA 'Charlie' that - like EVA 'Bravo' - is considered a "full" EVA. This means that all timeline content needs to be done in one EVA due to Node-2 Thermal Loop System shutdown timing. Also, working with ammonia (NH3) fluid couplings could require additional time for decontamination procedures.)

FE-1 Malenchenko completed the daily routine maintenance of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM, including ASU toilet facilities systems/replaceables.

Yuri's daily maintenance of the IMS (Inventory Management System), updating/editing its standard 'delta file', including locations, for the regular weekly automated export/import to its three databases on the ground (Houston, Moscow, Baikonur) today was a voluntary item on his 'time permitting' discretionary task list.

The FE-1 used the SKDS CMS (Pressure Control & Atmosphere Monitoring System/Countermeasure System) to take the periodic readings of potentially harmful contaminants in the SM. The hardware was then returned to initial stowage. (The CMS, part of the Russian GANK-4M atmospheric analysis suite, uses preprogrammed microchips to measure Formaldehyde (H2CO, methanal), Carbon Monoxide (CO) and Ammonia (NH3), taking one measurement per microchip.)

The crewmembers completed their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (CDR, FE-2), TVIS treadmill (FE-1), RED resistive exerciser (CDR, FE-2) and VELO bike with bungee cord load trainer (FE-1).

Afterwards, Dan transferred the crew's exercise data file to the MEC (Medical Equipment Computer) laptop for downlink, including the daily wristband HRM (Heart Rate Monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).

At ~3:35pm, the FE-2 is scheduled for another PFC (Private Family Conference) via S-band/audio and Ku-band/MS-NetMeeting application (which displays the uplinked ground video on the SSC-9 laptop).

Task items waiting for Whitson & Tani on their discretionary U.S. 'job jar' list include -

Work on BCAT-3 (Binary Colloidal Alloy Test-3, specifically remove samples from the CGBA-4 (Commercial Generic Bioprocessing Apparatus-4) & unstick the mixing magnet, with photography;
Relocating, in Node-2, the PS-120 (Power Strip) Junction Box from UOP-1 (Utility Outlet Panel 1) to the failed UOP-2; and
Starting U.S. trash gathering for disposal in Progress M-61/26P, aided by an uplinked preliminary (pre-E1) list of items.
EVA-12 'Charlie' Preview (Update 11/22): EVA-12 on Saturday begins nominally at ~5:00AM EST, probably much earlier real-time (knowing Peggy), with a 6h 30m timeline. Dedicated to Part 2 of Node-2 outfitting (i.e., connecting & configuring the second half of Harmony's fluid, power and cooling jumpers), its objectives are for Peggy Whitson & Dan Tani to:

Remove the S0 truss NH3 (ammonia) shunt jumper on Port;
Configure vent tools;
Remove Node-2 fluid QD (quick disconnect) caps;
Vent & stow the Port NH3 shunt jumper;
Relocate an APFR (Articulating Portable Foot Restraint) from Lab WIF-11 (Worksite Interface Fixture 11) to Lab WIF-12;
Relocate Node-2 Loop B fluid tray to the Lab & attach it;
Deploy Node-2 Loop B fluid tray hinged section;
Mate S0 fluid QDs, then open S0 valves and 2 fluid QDs;
Connect two Node-2 fluid line heater cables;
Connect five Node-2 Stbd avionics cables to Node-2;
Release Node-2 Stbd CBM (Common Berthing Mechanism) petal launch locks (8 latches);
Mate remaining SSPTS umbilical (connector J16A) to PMA-2 (connector; P16) (New)
Install Lab/Node-2 gap spanner on two handrails (Lab & Node 2); and (as get-ahead)
Release Node-2 Port CBM petal launch locks (8 latches). (Nadir CBM launch locks remain closed at this time to prevent the unlatched petals from opening (due to lost pin);
Remove Stbd SARJ (Solar Alpha Rotary Joint) Cover 7, using PGT (Pistol Grip Tool) on its 6 bolds.) (New)
Newly added Get-ahead tasks are:

Re-installation of CETA (Crew & Equipment Translation Aid) Light Fixture;
Clean-up of Lab MMOD (micrometeoroid/orbital debris) shield.

Node-2 Checkout Status: After crew ingress in Node-2 yesterday, ground controllers successfully executed Part 1 of Node-2 Activation & Checkout. Powered on & verified were Node-2 connections to the MTL (Moderate Temperature Loop), ETCS (External Thermal Control System) Loop A, and EPS (Electrical Power System) Channel 1/4. It has been determined that Node-2 will remain activated during EVA-12, although its interior lighting and the functioning UOP (Utility Outlet Panel) will be turned off. (Checkout of one string of the Node-2 EPS via activation of the DDCUs (DC-to-DC Converter Units) and RPCMs (Remote Power Control Modules) was successfully completed, as were the activations of the #1 MDM (Multiplexer/Demultiplexer) computer of Node-2 and of the Node-2 MTL. CSA Robotics successfully performed a brief SSRMS (Space Station Remote Manipulator System) base change to the Node-2 PDGF (Power and Data Grapple) to check out that connection, afterwards moving the robotarm back to the Lab PGDF to support EVA-12 viewing by its cameras. Node-2 Stbd & Port CBMs (Common Berthing Mechanisms) petals are properly latched so that the EVA-12 spacewalkers can remove their launch locks. The Nadir CBM, with petals unlatched because of the hatch pip pin lost in their mechanism, for the time being remains safed with launch locks engaged.)

CEO (Crew Earth Observation) photo targets uplinked for today were Mount Nyiragongo (this is one of Africa's most notable volcanic peaks and is located near the eastern border of the Democratic Republic of the Congo, just north of Lake Kivu, with much larger Lake Victoria well to the east. It is an active and dangerous volcano with steep slopes and extremely fluid lava flows. ISS pass was near nadir in mid-morning light. Frequent cloudiness makes this a challenging target, but the crew looked for landmarks and tried for long-lens, detailed views), Bosumtwi Impact Crater (this well-marked impact crater is located about 150 km west of the south end of Lake Volta in south central Ghana. It is a very young impact (just over a million years old), about 10.5 km in diameter, and almost completely filled by a lake. There are only a few images of this crater in the CEO database because the area is usually cloud and/or haze covered. On this early morning pass, as ISS approached the coast from the NW, the crew was to find Lake Volta and look just right of track), Kingman Reef, Central Pacific (Kingman Reef is a largely submerged, uninhabited tropical atoll located in the North Pacific Ocean, roughly half way between Hawaiian Islands and American Samoa. It is the northernmost of the Northern Line Islands and lies 65 km NNW of Palmyra Atoll, the next closest island. The total area within the rim of the reef, which has greater depths in the western part, is 60 km . There is just one small strip of dry land on the eastern rim, with an area of less than 0.01 km . The ISS pass was near nadir in early morning light, requiring long-lens settings for detail), and Palmyra Atoll, Central Pacific (Palmyra is one of the Northern Line Islands of the North Pacific Ocean and lies to the S-SE of Kingman Reef (previous CEO target). This oddly shaped atoll has an area of 4.6 square miles (12 km ), and it consists of an extensive reef, two shallow lagoons, and some 50 sand and reef-rock islets and bars covered with vegetation with a small air strip on the north side. ISS pass was near nadir in early morning, less than 10 seconds after Kingman).

Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
© 1997-2017 Mark Wade - Contact
© / Conditions for Use